radomes.openskiesproject.org
summerstormnight2002
summerstormnight2002 summerstormnight2002
  • 15-08-2016
  • Mathematics
Answered

Suppose you have the 26 letters of the alphabet on separate cards in a hat. You pick out a card, write down the letter, put the card back in the hat, mix up the cards, pick out another card, and so on. What is the probability that on your first four draws your letters spell MATH?

Verified Answer :

texaschic101
texaschic101 texaschic101
  • 15-08-2016

1/26  1/26 1/26 * 1/26 = 1/456976

Answer Link
samreid65 samreid65
  • 05-12-2018

Answer:


Step-by-step explanation:


Answer Link

Related Questions

Determine the mass of 305 moles of diatomic oxygen O2
How is the process of finding equivalent ratios like the process of finding equivalent fractions
The scientist who first studied the world of microscopic life was
What numbers can I multiply to get 24 and add to get negative 4
There was an Open Door policy between the US and the other countries that traded in China True False
when communication take place people often confuse sympathy and with empathy compare at least four difference between the two modes of communication
As soon as she got her console out of the package she pressed the power switch Within seconds the green lights appeared and smile crept across Jessicas face Thi
Intergrate secxsecxtanx dx
how do you simplify 37 4m
How did nationalism affect the Balkans